OpenStudy (anonymous):

Write each polar equation in rectangular form. θ=π/3 r=2cscθ

Parth (parthkohli):

If \((a,b)\) are the rectangular coordinates, then \((a,b) = (r\cos \theta ,r\sin\theta)\)

Parth (parthkohli):

Can you first figure \(r\) out?